# Molecular Dynamics and Machine Learning Unlock Possibilities in Beauty Design—A Perspective

### v1

@@ Abstract
While traditional fingerprint Wigh, Goodman, and Lapkin (2022)-based methods are still used, small molecules are now more ubiquitously modeled as graphs, and processed through a graph neural network (GNN) Kipf and Welling (2016a); Xu et al. (2018); Gilmer et al. (2017); Hamilton, Ying, and Leskovec (2017); Battaglia et al. (2018); Wang and Cho (2024) to form useful representations for downstream prediction tasks. Once the backbone of this mapping is constructed, it can also be used to quantify the uncertainty of its predictions in a Bayesian setting Wang and Karaletsos (2021); Wang et al. , for example with variational inference Blei, Kucukelbir, and McAuliffe (2017), or in an active learning setting to guide the design of the next round of experiments Kristiadi et al. (2024).
@@
In this perspective, we first review the overarching goals and specific challenges in beauty design. Next, we identify the physical (MD-based) and machine learning toolboxes useful in modeling and optimizing the propoerties of interest in beauty design. Finally, we also envision beauty design areas where MD simulation and ML modeling can be of immediate utility, and propose feasible projects.
@@ III.2 Micro hyaluronic acids: potency and risks
Micro hyaluronic acids with low molecular weight are well known for their ability to penetrate the skin effectively. These tiny molecules can go through the skins’ protective barrier and provide moisture directly to the deeper layers of the skin improving its elasticity and minimizing noticeable aging signs. This property is crucial in cosmetic dermatology, where hydration and skin suppleness are key to a youthful appearance. The effectiveness of micro hyaluronic acids in promoting skin hydration and reducing wrinkles has been substantiated by various studies, which highlight its efficacy as a humectant, with the ability to hold water up to a thousand times their weight and facilitate better absorption compared to their higher molecular counterparts Bravo et al. (2022); Juhaščik, Kováčik, and Huerta-Ángeles (2022). Advancements in skincare technology have leveraged the properties of micro hyaluronic acids to develop sophisticated delivery systems that maximize skin benefits. Techniques such as encapsulation in liposomes or binding to nanoparticles have been explored to further enhance the penetration and longevity of hyaluronic acids in skin layers. These innovations not only improve the immediate hydrating effects but also contribute to the long-term anti-aging benefits by maintaining higher levels of skin hydration over extended periods Juhaščik, Kováčik, and Huerta-Ángeles (2022).
@@
Characterizing Equation 1 by drawing samples while integrating Newton’s second motion of law, MD simulation finds use in all schools of molecular modeling. There have been a wide variety of mature and wide used pipelines for modeling the energy landscape of biomolecular systems of therapeutical interest, which can be seamlessly adopted to model those of cosmeceutical significance.
@@
For beauty design, MD simulations Case et al. (2021); Lee et al. (2016); Abraham et al. (2015); Eastman et al. (2023) can be employed to model the interactions between the cosmeceutical molecules and their targets as well as provide detailed insights into the hydrogen bonding interactions that are critical for binding specificity and stability Decherchi and Cavalli (2020); Agu et al. (2023). Quantitatively, the binding free energy between the small molecule and the target protein can be reliably predicted Mobley and Gilson (2017). For crude predictions, popular methods include MM-GBSA (Molecular Mechanics Generalized Born Surface Area) and MM-PBSA (Molecular Mechanics Poisson-Boltzmann Surface Area) Wang et al. (2019); Genheden and Ryde (2015). Alanine scanning methods Liu et al. (2018); Yan et al. (2017) can be used to further deduce the contribution of each residue towards the binding free energy. More elaborately, free energy calculations based on thermodynamic integration (TI) Bruckner and Boresch (2011) and free energy perturbation (FEP) Jorgensen and Thomas (2008); Mey et al. (2020); Shirts, Mobley, and Chodera (2007) have shown best correlation with experimental data, and therefore can be used to reliably predict the binding free energy between the small cosmeceutical molecule and the target.
Physical property prediction
@@
Usually, Cosmetic compounds necessitate a thorough evaluation of ADMET profiles to guarantee their suitability for topical application Martins et al. (2014). This comprehensive assessment ensures that active ingredients are effectively absorbed through the skin barrier, distributed appropriately within the skin layers, metabolized without generating harmful by-products, and excreted without accumulation that could lead to adverse effects. Traditionally, ADMET evaluation has relied primarily on experimental methods to determine these key pharmacokinetic and toxicological properties of drugs or active molecules. Recently, the application of machine learning (ML) and deep learning (DL) techniques in ADMET prediction has emerged as a complementary approach to experimental methods, offering advantages in terms of speed and accuracy Cronin et al. (2022); Xin et al. (2024); Wilm, Kühnl, and Kirchmair (2018). Besides, Physical property prediction is required in the product’s fluid properties when applied to the skin. These properties include viscosity, rheological behavior, surface tension, diffusion coefficients, and absorption kinetics Lochhead et al. (2017); Lochhead (2017). These factors collectively determine a cosmetic product’s ease of application, spreadability, and stability. Advanced analytical techniques such as atomic force microscopy and nuclear magnetic resonance spectroscopy provide molecular-level structural and dynamic information, further aiding in the understanding of how molecular interactions influence macroscopic physical properties. Employing predictive methods for the design of cosmetic products offers a cheap method to predict the cosmetic molecule’s behavior. Computational method such as computational fluid dynamics, molecular simulations, machine learning enable the forecasting of a product’s behavior under various conditions, facilitating the optimization of formulations before physical prototypes are developed.
Spectroscopy prediction
During the product development stage, spectroscopic techniques such as near-infrared (NIR), Raman, and UV-visible spectroscopy are employed to characterize raw materials and analyze formulations Paudel, Raijada, and Rantanen (2015). Spectroscopic prediction serves as a valuable alternative to costly experimental procedures in the cosmetics industry, as it allows the prediction of product performance and properties using existing spectral data and mathematical models, either before or during experimentation. By analyzing the spectral characteristics of raw materials, spectroscopic prediction can anticipate their behavior and efficacy within a formulation. Drawing on physical theories regarding how molecular structures influence macroscopic properties, researchers can logically deduce the interactions between different components, thereby rapidly identifying the most suitable ingredients Kowalczuk et al. (2021). For instance, by predicting the spectral properties of blended components and leveraging the physical characteristics of light absorption, reflection, and scattering, it is possible to optimize the formulations of products such as lipsticks and foundations, which require precise color differentiation Rico et al. (2023). Moreover, spectroscopic prediction can be utilized to forecast product stability Prieto et al. (2009). For example, in sunscreen products, spectroscopic prediction models can quickly estimate SPF values by leveraging the physical relationship between UV absorption and sunscreen efficacy, thereby reducing the need for extensive in vitro or in vivo testing Cole, Silverman, and Bonitatibus (2019).
@@ VI The machine learning toolbox
where $\phi,\rho$ are the update and aggregate function, repsectively. Omitting the nonlinear transformation step $\phi$ ubiquitous in all neural models, and assuming a convolutional aggregate function, $\rho=\operatorname{SUM}$ or $\rho=\operatorname{MEAN}$, a graph neural network layer is characterized by the aggregation/convolution operation that pools representations from neighboring nodes, forming an intermediary representation $\mathbf{X}^{\prime}$, which on a global level, with activation function $\sigma$ and weights $W$, can be written as: $\mathbf{X}^{\prime}=\sigma(\hat{A}\mathbf{X}W)$. The primary difference among architectures amounts to distinct effective adjacency matrix $\hat{A}$. The most classical examples include: graph convolutional networks Kipf and Welling (2016a) (GCN), which normalizes $\hat{A}$ by the in-degree of nodes $D_{ii}=\sum_{j}A_{ij}$ and graph attention networks velickovic2018graph (GAT), which takes $\hat{A}$ as the attention score;
@@
Once we have a reliable model for the prediction of activities given the structure of the graph, we can couple it with a generative model for the true de novo design of molecules and pipetides. This can be achieved through a principled, probabilistic manner, where the conditional distribution in Equation 4 is replaced by the distribution of the chemical graphs without measurements:
@@
With some over-simplification, molecular discovery can be abstracted as a sequential active learning process. Given a finite pool of candidate chemical compounds, represented by chemical graphs =$\mathcal{F}=\{\mathcal{G}_{i},i=1,2,\dots N\}$, each associated with a potency $y(\mathcal{G}_{i})\in\mathbb{R}$, of which a noisy observation is expensive to acquire, we are interested in finding the compound with the highest potency $\mathcal{G}^{*}=\operatorname{argmax}_{i}y(\mathcal{G}_{i})$ with fewest function evaulations. This can be achieved via active learning, starting with an empty portfolio $\mathcal{P}=\emptyset$, and in each round, choosing a compound from the candidate pool, $\mathcal{F}$, based on a decision informed by the current portfolio, $\hat{\mathcal{G}}=D(\mathcal{F}|\mathcal{P}),\hat{\mathcal{G}}\in\mathcal{F}$. Subsequently, we subtract this compound from the candidate pool to add it to the portfolio, $\mathcal{F}\leftarrow\mathcal{F}\setminus\{\hat{\mathcal{G}}\},\mathcal{P}\leftarrow\mathcal{P}\cup\{\hat{\mathcal{G}}\}$.
@@ VII Roadmap: Potential use cases
Inhibiting the activity of tyrosinase, an essential enzyme for melanin synthesis in skin, arbutin has been ubiquitously used as skin-lightening agents targeting the hyperpigmentation issues. Usually extracted from the leaves of plants, such as the bearberry, the efficacy of this natural product is bottlenecked by its inability to penetrate skin to reach the melanocytes Masyita et al. (2021). We propose to develop arbutin derivatives displaying similar efficacy while showing significantly greater lipophilicity for better skin penetration.
@@
The extremely small molecular weight of hyluronic acid is both a blessing and a curse—it enables the easy penetration of the skin while also triggers inflammatory reactions Bravo et al. (2022). The chemical modification of hyluronic acid Bokatyi, Dubashynskaya, and Skorik (2024) has been proven to be an effective avenue towards efficient drug delivery. Using the modification strategies outlined in Bokatyi, Dubashynskaya, and Skorik (2024), we can chemically decorate the backbone with inactive functional groups to moderately increase the molecular weight of hyluronic acid, while not altering its ability to hold water. These modifications can be propsoed either by generative models or fragment-based structure enumeration and library generation tools such as Polishchuk (2020), while the physical properties and the ADMET profile can be predicted using a machine learning-enabled QSAR model.
@@
The majority of peptide design in the current beauty and skincare market still relies heavily on traditional, resource-intensive methods that are both time-consuming and laborious. These conventional approaches often result in peptides that lack the necessary modifications to enhance their skin penetration and bioactivity. However, with the rapid advancements in deep learning, particularly the application of diffusion models, peptide sequence prediction and optimization have become more efficient and precise. By training diffusion models on extensive datasets of known cosmeceutical peptides, researchers can capture the complex patterns in their sequence distribution and functional properties, enabling the generation of novel peptides with desired characteristics. Specifically, advanced peptide design using diffusion models can utilize existing cosmetic peptides as input and, through conditional generation, produce new sequences that exhibit multiple skincare benefits. For instance, these generated peptides can be designed to possess not only antioxidant and anti-inflammatory properties but also promote collagen synthesis and provide skin-whitening effects. Moreover, through multiple iterations of model evaluation and screening, the efficacy and safety of the generated peptides can be validated, ensuring their reliability in practical applications. During the peptide design process, targeted chemical modifications can be incorporated, such as cyclization, lipid chain conjugation, or glycosylation of terminal groups, to enhance the peptides’ penetration ability and stability. These modifications not only improve the efficiency of peptide translocation across the skin barrier but also increase their bioactivity at the target site, thereby ensuring the effective realization of multiple cosmetic benefits. Through iterative evaluation and selection, researchers can continuously optimize and identify the most promising peptide sequences, eliminating ineffective or potentially harmful molecules, thus significantly enhancing the quality and efficacy of the final products.

### v2

@@ Abstract
While traditional fingerprint-based Wigh, Goodman, and Lapkin (2022) methods are still used, small molecules are now more ubiquitously modeled as graphs, and processed through a graph neural network (GNN) Kipf and Welling (2016a); Xu et al. (2018); Gilmer et al. (2017); Hamilton, Ying, and Leskovec (2017); Battaglia et al. (2018); Wang and Cho (2024) to form useful representations for downstream prediction tasks. Once the backbone of this mapping is constructed, it can also be used to quantify the uncertainty of its predictions in a Bayesian setting Wang and Karaletsos (2021); Wang et al. , for example with variational inference Blei, Kucukelbir, and McAuliffe (2017), or in an active learning setting to guide the design of the next round of experiments Kristiadi et al. (2024).
@@
In this perspective, we first review the overarching goals and specific challenges in beauty design. Next, we identify the physical (MD-based) and machine learning toolboxes useful in modeling and optimizing the properties of interest in beauty design. Finally, we also envision beauty design areas where MD simulation and ML modeling can be of immediate utility, and propose feasible projects.
@@ III.2 Micro hyaluronic acids: potency and risks
Micro hyaluronic acids with low molecular weight are well known for their ability to penetrate the skin effectively. These tiny molecules can go through the skin’s protective barrier and provide moisture directly to the deeper layers of the skin improving its elasticity and minimizing noticeable aging signs. This property is crucial in cosmetic dermatology, where hydration and skin suppleness are key to a youthful appearance. The effectiveness of micro hyaluronic acids in promoting skin hydration and reducing wrinkles has been substantiated by various studies, which highlight its efficacy as a humectant, with the ability to hold water up to a thousand times their weight and facilitate better absorption compared to their higher molecular counterparts Bravo et al. (2022); Juhaščik, Kováčik, and Huerta-Ángeles (2022). Advancements in skincare technology have leveraged the properties of micro hyaluronic acids to develop sophisticated delivery systems that maximize skin benefits. Techniques such as encapsulation in liposomes or binding to nanoparticles have been explored to further enhance the penetration and longevity of hyaluronic acids in skin layers. These innovations not only improve the immediate hydrating effects but also contribute to the long-term anti-aging benefits by maintaining higher levels of skin hydration over extended periods Juhaščik, Kováčik, and Huerta-Ángeles (2022).
@@
Characterizing Equation 1 by drawing samples while integrating Newton’s second motion of law, MD simulation finds use in all schools of molecular modeling. There have been a wide variety of mature and widely used pipelines for modeling the energy landscape of biomolecular systems of therapeutical interest, which can be seamlessly adopted to model those of cosmeceutical significance.
@@
To design effective cosmetic molecules, molecular dynamics (MD) simulations Case et al. (2021); Lee et al. (2016); Abraham et al. (2015); Eastman et al. (2023) can be employed to model the interactions between cosmeceutical molecules and their biological targets. To be more specific, MD simulations enable the accurate prediction of binding free energy, a key metric in assessing the strength and stability of the interaction between a small molecule, such as a cosmetic agent, and a target protein Mobley and Gilson (2017). One of these methods is commonly categorized as endpoint techniques, such as molecular mechanics generalized Born surface area (MM-GBSA) and molecular mechanics Poisson-Boltzmann surface area (MM-PBSA), to estimate free energy changes by evaluating the energy differences between the bound and unbound states with data sampled from the same molecular dynamics trajectory Wang et al. (2019); Genheden and Ryde (2015). In both approaches, the binding free energy is divided into several components, including enthalpic and entropic contributions. The former component is calculated as the sum of the gas-phase interaction energy between the ligand and protein and the solvation free energy. At the same time, the latter is computed as conformational entropy change related to ligand binding Tuccinardi (2021). To enhance computational efficiency, these endpoint methods employ an implicit continuum solvent model, which significantly reduces the complexity of the calculations by omitting explicit water molecules from the MD simulations. While this approach accelerates calculations, it may also compromise the accuracy, as it overlooks the potentially significant role of specific water-mediated interactions. Besides, a key limitation of MM-PBSA and MM-GBSA is their treatment of entropy. These methods often neglect entropic contributions or handle them in a less rigorous manner. For instance, the configurational entropy is either approximated using simplified models or ignored altogether Wang et al. (2016). Therefore, MM-GBSA/PBSA exhibit moderate accuracy while being computationally efficient, which makes them well-suited for the rapid screening of potential small molecules in the early stages of drug discovery, such as anti-tumor and anti-virus drugs Hu and Contini (2019); Forouzesh and Mishra (2021); Shridhar Deshpande et al. (2021). For the calculation of complex interactions and highly flexible ligands, we developed the alanine scanning with generalized Born and interaction entropy (ASGBIE) method Liu et al. (2018); Yan et al. (2017). Compared to MM-GBSA, ASGBIE provides more accurate binding free energy estimations derived from improved solvation energy calculations and a more rigorous consideration of entropic contributions, which are often neglected in MM-GBSA. Additionally, the method employs variable dielectric coefficients ranging from 1 to 10 for different categories of amino acids, including non-polar, polar, and charged groups, allowing for more nuanced electrostatic calculations. In a comparison of ASGBIE and MM-GBSA on 24 complex systems with 433 pairs of mutation, ASGBIE achieved a mean unsigned error (MUE) of 0.2 kcal/mol, outperformed the MUE of 1.0 kcal/mol for MM-GBSA Yan et al. (2017). Additionally, ASGBIE has been applied to predict hotspots with the highest contributions in binding free energy on a protein-peptide system, aiding peptide modifications to enhance binding affinity, and estimating relative binding free energy changes between wild-type and mutated peptides with reported accuracy Zhou et al. (2024).
More sophisticated methods can also be considered for estimating binding free energies. One important category is alchemical methods, including thermodynamic integration (TI) Kirkwood (1935) and free energy perturbation (FEP) Zwanzig (1954). These methods estimate free energy changes by sampling intermediate states between initial and final configurations based on sampling configurations guided by the potential energy of the initial state. Two more extended functions were then developed based on the FEP equation. Bennett derived a Bennett acceptance ratio (BAR) Bennett (1976) function requiring sampling configuration at both states to minimize the expected square error of the free energy difference. Meanwhile, a multistate Bennett acceptance ratio (MBAR) Shirts and Chodera (2008) function takes into account simulation data from multiple states, involving a more comprehensive sampling. These methods involve the calculation of free energy difference between two states by introducing a coupling parameter ($\lambda$) that varies from 0 to 1 to sample configurations for both states, either based on the ensemble average of the state 0 or that of both states, creating a series of intermediate states that improve phase space overlap and facilitate a smoother transformation. The utilization of soft-core potential further avoids the dispersed energy distribution on endpoints of 0 and 1. Furthermore, alchemical methods have been extensively applied in estimating binding free energy between target proteins and small molecules, such as large-scale screenings or docking poses evaluation Bhati and Coveney (2022); Goldfeld et al. (2015); Kaus et al. (2015). Some approaches combine alchemical methods with replica exchange solute tempering (FEP/REST), further improving accuracy by overcoming energy barriers associated with local minima Bourion-Bédès et al. (2021); Lockhart et al. (2023); Ngo et al. (2020). Several software platforms support alchemical methods. For instance, AMBER18 enables GPU-accelerated TI calculations, and AMBER20 package was evaluated for its accuracy on a dataset of around 200 ligand mutations spanning 8 protein targets, achieving a mean unsigned error (MUE) of 0.84 kcal/mol Lee et al. (2020). Similarly, a commercial software FEP+ exploited FEP/REST approach on the same dataset and achieved a comparable MUE of 0.86 kcal/mol Wang et al. (2015). These alchemical approaches offer more efficient sampling of the protein-ligand phase space and have demonstrated strong correlations with experimental data, making them reliable for accurately predicting the binding free energy of small cosmeceutical molecules interacting with their targets. Another category of methods includes enhanced sampling techniques like umbrella sampling and metadynamics, which are useful for large-scale molecular systems Williams-Noonan, Yuriev, and Chalmers (2018). However, these methods significantly increase computational demand due to the exponential growth in the number of molecules and interactions, necessitating the use of high-performance computing (HPC) resources and substantial storage capacity.
Physical property prediction.
@@
Usually, Cosmetic compounds necessitate a thorough evaluation of ADMET profiles to guarantee their suitability for topical application Martins et al. (2014). This comprehensive assessment ensures that active ingredients are effectively absorbed through the skin barrier, distributed appropriately within the skin layers, metabolized without generating harmful by-products, and excreted without accumulation that could lead to adverse effects. Traditionally, ADMET evaluation has relied primarily on experimental methods to determine these key pharmacokinetic and toxicological properties of drugs or active molecules. Recently, MD simulations have emerged as an important tool for aiding in the computational analysis of ADMET Srivastava et al. (2025); Balkrishna et al. (2024). By simulating the behavior of molecules within cell membranes and skin tissues, researchers can predict how these molecules penetrate the skin barrier, their distribution within the skin, and their interactions with proteins. Additionally, MD simulations can be combined with other computational methods, such as machine learning and computational fluid dynamics, to predict the fluid properties of products under various conditions, including viscosity, surface tension, and diffusion coefficients. These predictions can effectively guide the design of cosmetic formulations, reduce development costs, and improve the application performance and stability of products.
@@ VI The machine learning toolbox
where $\phi,\rho$ are the update and aggregate function, repsectively. Omitting the nonlinear transformation step $\phi$ ubiquitous in all neural models, and assuming a convolutional aggregate function, $\rho=\operatorname{SUM}$ or $\rho=\operatorname{MEAN}$, a graph neural network layer is characterized by the aggregation/convolution operation that pools representations from neighboring nodes, forming an intermediary representation $\mathbf{X}^{\prime}$, which on a global level, with activation function $\sigma$ and weights $W$, can be written as: $\mathbf{X}^{\prime}=\sigma(\hat{A}\mathbf{X}W)$. The primary difference among architectures amounts to distinct effective adjacency matrix $\hat{A}$. The most classical examples include: graph convolutional networks (GCN) Kipf and Welling (2016a), which normalizes $\hat{A}$ by the in-degree of nodes $D_{ii}=\sum_{j}A_{ij}$ and graph attention networks (GAT) Veličković et al. (2017), which takes $\hat{A}$ as the attention score;
@@
Once we have a reliable model for the prediction of activities given the structure of the graph, we can couple it with a generative model for the true de novo design of molecules and peptides. This can be achieved through a principled, probabilistic manner, where the conditional distribution in Equation 4 is replaced by the distribution of the chemical graphs without measurements:
@@
With some over-simplification, molecular discovery can be abstracted as a sequential active learning process. Given a finite pool of candidate chemical compounds, represented by chemical graphs =$\mathcal{F}=\{\mathcal{G}_{i},i=1,2,\dots N\}$, each associated with a potency $y(\mathcal{G}_{i})\in\mathbb{R}$, of which a noisy observation is expensive to acquire, we are interested in finding the compound with the highest potency $\mathcal{G}^{*}=\operatorname{argmax}_{i}y(\mathcal{G}_{i})$ with fewest function evaluations. This can be achieved via active learning, starting with an empty portfolio $\mathcal{P}=\emptyset$, and in each round, choosing a compound from the candidate pool, $\mathcal{F}$, based on a decision informed by the current portfolio, $\hat{\mathcal{G}}=D(\mathcal{F}|\mathcal{P}),\hat{\mathcal{G}}\in\mathcal{F}$. Subsequently, we subtract this compound from the candidate pool to add it to the portfolio, $\mathcal{F}\leftarrow\mathcal{F}\setminus\{\hat{\mathcal{G}}\},\mathcal{P}\leftarrow\mathcal{P}\cup\{\hat{\mathcal{G}}\}$.
@@ VII Roadmap: Potential use cases
Inhibiting the activity of tyrosinase, an essential enzyme for melanin synthesis in skin, arbutin has been ubiquitously used as skin-lightening agents targeting hyperpigmentation issues. Usually extracted from the leaves of plants, such as the bearberry, the efficacy of this natural product is bottlenecked by its inability to penetrate the skin to reach the melanocytes Masyita et al. (2021). We propose to develop arbutin derivatives displaying similar efficacy while showing significantly greater lipophilicity for better skin penetration.
@@
The extremely small molecular weight of hyaluronic acid is both a blessing and a curse—it enables the easy penetration of the skin while also triggering inflammatory reactions Bravo et al. (2022). The chemical modification of hyaluronic acid Bokatyi, Dubashynskaya, and Skorik (2024) has been proven to be an effective avenue towards efficient drug delivery. Using the modification strategies outlined in Bokatyi, Dubashynskaya, and Skorik (2024), we can chemically decorate the backbone with inactive functional groups to moderately increase the molecular weight of hyaluronic acid, while not altering its ability to hold water. These modifications can be proposed either by generative models or fragment-based structure enumeration and library generation tools such as Polishchuk (2020), while the physical properties and the ADMET profile can be predicted using a machine learning-enabled QSAR model.
@@
Peptides are currently predominantly composed of oligopeptides (containing fewer than 10 amino acids), forming a vast polypeptide space. The design of synthetic peptides still relies heavily on traditional, resource-intensive methods that are both time-consuming and laborious. These conventional approaches often result in peptides that lack the necessary modifications to enhance their skin penetration and bioactivity. However, advancements in technology have made peptide sequence prediction and optimization more efficient and precise. In terms of de novo peptide design, researchers can now leverage diffusion models trained on medicinal peptide datasets to capture the complex patterns of peptide sequence distribution and functional properties, enabling the generation of novel peptides with desired characteristics. Additionally, protein function prediction tools, such as DeepFRI, can be employed to predict the corresponding protein functions of these peptides Gligorijević et al. (2021). From a peptide modification perspective, targeted chemical modifications can be introduced to enhance the permeability and stability of peptides. For instance, acetyl tetrapeptide-15, a chemically modified version of Tyr-Pro-Phe-Phe, demonstrates how modifications like cyclization, lipid chain conjugation, or terminal glycosylation can be applied. These modified peptides can then be evaluated using molecular dynamics (MD) simulations. Through iterative cycles of model evaluation and screening, the efficacy and safety of the generated peptides can be validated.